OpenStudy (anonymous):

Davida bought a pair of earrings, each in the shape of an isosceles triangle with equal side lengths of 24 mm and a base length of 16 mm. She used one earring to make another pair of earrings, similar in shape, with the base length equal to 11 mm. What is the length of the equal-length sides of this new pair of earrings? mm

OpenStudy (anonymous):

2 medal rewarc

OpenStudy (anonymous):

reward

OpenStudy (anonymous):

what can you guess @urbanmorgans

OpenStudy (anonymous):

384

OpenStudy (anonymous):

i times the two #s

OpenStudy (agl202):

new side/24 = newbase/oldbase newside/24 = 11/16 newside = 11(24)/16 = 16.5 mm

OpenStudy (agl202):

16.5 mm is correct :)

OpenStudy (anonymous):

can u help me on another one

OpenStudy (agl202):

Alright, so were is my reward?